# Heterogeneity in Entity Matching: A Survey and Experimental Analysis

### v1

@@ Abstract
This survey offers a unified perspective on HEM by introducing a taxonomy, grounded in prior work, that distinguishes two primary categories—representation and semantic heterogeneity—and their subtypes. The taxonomy provides a systematic lens for understanding how variations in data form and meaning shape the complexity of matching tasks. We then connect this framework to the FAIR principlesFindability, Accessibility, Interoperability, and Reusability—demonstrating how they both reveal the challenges of HEM and suggest strategies for mitigating them.
@@
This is a preprint. The final version may be available via the publisher upon acceptance.
Keywords: Entity Matching, Entity Resolution, Data Heterogeneity
@@ 1 Introduction
Entity Matching (EM) has long been a fundamental component of data integration, cleaning, and analytics pipelines. Although recent advances—especially in deep learning and AI—have accelerated progress [11, 16, 41], EM systems continue to struggle in real-world deployments. High-performing models trained on clean, benchmark datasets struggle to generalize when exposed to messy, noisy, and heterogeneous data found in practice. These failures are not incidental; they stem from a pervasive and under-addressed challenge: data heterogeneity. Differences in formats (e.g., dates, units), schemas (e.g., attribute names, nesting), terminology (e.g., synonyms, language), and data quality (e.g., missing or inconsistent values) introduce mismatches in structure, semantics, and quality between development and deployment settings. Such heterogeneity undermines blocking, feature extraction, and similarity computation, leading to degraded performance across the entire EM pipeline. Even recent deep learning-based methods, which perform well on standard benchmarks, suffer sharp drops in accuracy when applied across domains with varying schema or semantics [50, 39].
@@
While heterogeneity in data is well-recognized across many domains—including information retrieval [69], geospatial systems [27], the Internet of Things [2], and big data analytics [6]—it poses particularly acute and evolving challenges in EM. Historically, heterogeneity in EM has been handled under themes such as schema matching [63], duplicate record detection [20], and semantic integration [14], typically focusing on specific dimensions like attribute alignment mismatches or lexical variation. However, the modern data landscape—characterized by large-scale, semi-structured and unstructured sources from the web, data lakes, IoT devices, and enterprise systems—introduces more complex and compounded forms of heterogeneity. These include variation in data formats (e.g., JSON, XML, relational), schemas, semantics, language, granularity, and data quality. In this context, mismatches in data models and semantic assumptions severely complicate schema alignment, feature extraction, and record linkage [5]. These challenges call for EM methods that are explicitly robust to heterogeneity, and for a systematic understanding of the many forms that heterogeneity can now take in practice.
@@
To address this challenge, we argue that categorizing and systematically studying data heterogeneity is essential for advancing the design, evaluation, and deployment of EM systems. A principled taxonomy of heterogeneity is useful for organizing prior work and also enables several concrete benefits in the design and assessment of EM systems. First, it guides the development of targeted model architectures, allowing practitioners to align method design with the expected types of heterogeneity. Prior work has demonstrated that different model families excel under specific types of heterogeneity—for example, transformer-based architectures have shown robustness to semantic variation such as synonyms and abbreviations [17, 9, 73], while graph-based methods are effective at capturing structural mismatches across schemas [87, 78, 72, 36]. Building on this foundation, our paper provides new experimental evidence supporting the need for heterogeneity-aware modeling.
@@
Second, such a taxonomy enables component-level stress testing, in which researchers can evaluate how EM methods respond to controlled semantic or structural mismatch at different stages of the pipeline—from blocking to similarity computation to final classification. Third, it supports the construction of heterogeneity-aware benchmarks and perturbation-based evaluation frameworks. Finally, an explicit understanding of heterogeneity contributes to uncertainty quantification and robustness analysis, which are increasingly critical for deploying EM systems in high-stakes domains such as healthcare, scientific data integration, and finance.
@@
One of the most significant recent shifts in the EM landscape is the increasing influence of large language models (LLMs) and generative AI. These models offer new capabilities that are particularly relevant for addressing semantic heterogeneity, a core challenge in modern EM. Pretrained models such as BERT and GPT have demonstrated strong abilities to capture lexical and contextual variation through transfer learning, reducing the need for hand-crafted features or schema-specific engineering [39]. More recently, prompting and instruction tuning have enabled the use of foundation models for zero- or few-shot entity resolution [19], making it possible to generalize across domains without extensive retraining. These trends suggest that foundation models are poised to play a growing role in heterogeneity-aware EM—a theme we revisit in detail later in the survey.
The study of heterogeneity in EM also has significant implications for data governance and interoperability, particularly in the context of the FAIR principles for scientific data management—ensuring that data is Findable, Accessible, Interoperable, and Reusable [84]. Heterogeneity presents direct challenges to achieving FAIR compliance, especially when integrating records across fragmented, inconsistent, or mismatched sources. Conversely, EM methods that are explicitly designed to handle such heterogeneity can act as critical enablers of FAIRification by enhancing schema alignment, disambiguation, and record linkage. Throughout this paper, we emphasize the mutual relationship between EM and FAIR, and show how heterogeneity-aware EM systems contribute to building more trustworthy, transparent, and reusable data infrastructures.
@@
  • We introduce a hierarchical taxonomy of data heterogeneity in EM, distinguishing between representation heterogeneity (e.g., format, schema) and semantic heterogeneity (e.g., language, granularity, quality). This taxonomy serves as a conceptual tool for analyzing where and how EM systems break down in heterogeneous settings.
  • We systematically survey and categorize recent EM methods through the lens of this taxonomy, revealing how different model classes—rule-based, neural, and graph-based—address (or fail to address) specific forms of heterogeneity, and identifying patterns and gaps in current research.
@@ 2 A Framework for Classifying HEM
[…]
(Schema)] ] [Semantic [Terminology &
Language] [Contextual
Variability] [Granularity &
Resolution] [Temporal
Variability] [Data
Quality] ] ]
@@ 2.1 Representation Heterogeneity
Representation heterogeneity encompasses the structural and syntactic differences that occur when datasets use different modalities, formats, or schema designs to describe entities. These differences can disrupt every stage of the EM pipeline—from feature extraction to similarity computation—by introducing incompatibilities in how records are organized or encoded. For example, one dataset might store product data as nested JSON objects, while another uses flat CSV files. Even when datasets describe the same entities, format mismatches and inconsistent attribute organization can hinder alignment. Addressing representation heterogeneity typically requires schema matching, format normalization, or modality-specific processing techniques. We break this category into three subtypes:
  • Multimodality: Datasets may include diverse data types—such as text, images, and videos—for describing entities. For instance, e-commerce records may pair textual descriptions with product photos. Aligning entities across modalities requires models that can jointly embed or compare representations across heterogeneous data sources.
@@ 2.2 Semantic Heterogeneity
Early database work recognized these challenges in federated systems [67], and Semantic Web research later addressed ontology alignment [21]. Today, modern EM methods incorporate contextual modeling, domain adaptation, and knowledge resources to address semantic heterogeneity—topics we explore in depth in Section 3.
@@ 3 Review of HEM and Related Research Areas
To systematically review existing EM methods in relation to heterogeneity, we adopted a multi-step process. We first identified a set of key research areas that are central to managing heterogeneity in EM. These include schema matching and structural heterogeneity, representation learning and semantic embeddings, deep learning and graph neural networks, knowledge bases, transfer learning and domain adaptation, active learning, data lakes, and LLMs. These topics were selected not only for their relevance to HEM but also for their prominence in recent EM research. We then collected and reviewed articles published since 2019 that address EM in the context of any of these areas. Our selection focused on peer-reviewed papers from leading venues in data management, artificial intelligence, and machine learning. We excluded theses, posters, and articles that did not explicitly engage with heterogeneity. From an initial pool of 122 articles, we identified 31 core studies that directly address HEM, with particular emphasis on semantic heterogeneity and its intersections with structural and format heterogeneity, given their frequent overlap in practical EM settings.
3.1 Schema Matching and Structural Heterogeneity
Schema matching is a longstanding focus in data integration and is widely used in EM to address format and structural heterogeneity [63]. Recent research has expanded this view, introducing new benchmarks and algorithms to better handle heterogeneous schemas in modern EM settings.
One such contribution is Machamp [79], which introduces a benchmark for Generalized Entity Matching (GEM). Unlike earlier EM benchmarks that assume structured data with aligned schemas, Machamp includes tasks involving structured, semi-structured (e.g., JSON), and unstructured (e.g., text) data. It covers seven real-world scenarios that reflect schema mismatches, such as matching relational with semi-structured or textual data. By repurposing and transforming existing datasets, Machamp evaluates model robustness under schema and semantic variation. Experiments reveal substantial performance drops for deep models like BERT and DITTO in heterogeneous settings, underscoring the need for more schema-agnostic techniques.
@@ 3.1 Schema Matching and Structural Heterogeneity
Addressing this need, HERA [42] proposes a different paradigm for entity resolution. Instead of first aligning schemas, HERA operates directly on heterogeneous records, reducing information loss from premature schema integration. It uses a novel compare-and-merge process to iteratively build “super records,” combining both instance-based and schema-based similarity signals. An indexing structure supports efficient candidate generation and similarity computation. Empirical results show that HERA significantly outperforms state-of-the-art methods, particularly when schema heterogeneity is high and schema mapping is lossy.
@@
GraphER [36] also avoids explicit schema alignment by adopting a token-centric model built on Graph Convolutional Networks (GCNs). It constructs an Entity Record Graph (ER-Graph) that encodes relationships between records, attributes, and tokens. Through a two-layer GCN, GraphER jointly learns structural and semantic token embeddings, enabling fine-grained token-level comparisons without fixed attribute alignment. Evaluations on standard benchmarks show that GraphER consistently outperforms baseline models, especially in cases with diverse or sparse schemas.
@@ 3.2 Representation Learning and Semantic Embeddings
Representation learning has become foundational to modern EM, especially with the rise of deep neural models. Unlike static feature extraction, representation learning automatically identifies latent patterns that are most predictive for matching, enabling comparisons across diverse formats, schemas, and domains. Among these techniques, semantic embeddings—such as Word2Vec, GloVe, and BERT—encode the contextual meaning of tokens, attributes, and records. This makes them especially effective for resolving semantic heterogeneity, such as synonymy, polysemy, and language variation. Most recent EM systems use some form of learned representations to handle heterogeneity in both structure and meaning.
A leading example is DITTO [40], which fine-tunes Transformer-based models such as BERT and RoBERTa for EM tasks. Framing EM as a sequence-pair classification problem, DITTO uses contextual embeddings to capture language structure and relational cues. Its performance is further improved by (1) domain knowledge injection to highlight key tokens, (2) TF-IDF summarization to handle long records, and (3) data augmentation to enhance robustness. These design choices reduce dependency on large labeled datasets and allow the model to generalize across schemas without requiring attribute alignment.
Extending this line of work, HierGAT [87] introduces a Hierarchical Graph Attention Transformer that models entities, attributes, and tokens through a layered graph structure. It uses a combination of self-attention and graph attention to learn multi-level contextual embeddings. The model addresses challenges like polysemous terms, attribute salience, and noisy input by capturing both semantic and relational dependencies. HierGAT demonstrates strong performance on “dirty” datasets, showing robustness to data quality issues.
Beyond direct EM methods, several systems advance representation learning for heterogeneous data integration. Termite [22] learns a shared embedding space over structured and unstructured data, embedding rows, columns, and text snippets into a unified latent space. Its Termite-Join operator enables cross-format querying without schema alignment. Similarly, EmbDI [10] constructs embeddings using a tripartite graph over tokens, attributes, and tuples. Through random walks, it generates a corpus for training embeddings that encode both structural and semantic relationships, supporting robust matching even with noisy or sparse input.
@@
Complementing these architectures, [89] provide an empirical comparison of twelve off-the-shelf embeddings, including FastText, SBERT, and various BERT variants, across blocking and matching tasks. Surprisingly, they find that cosine similarity over frozen embeddings often rivals fine-tuned deep models, depending on dataset properties. These insights help practitioners select efficient and effective embedding strategies for heterogeneous matching scenarios.
EM Transformer [9] evaluates four transformer architectures—BERT, RoBERTa, XLNet, and DistilBERT—on noisy and textual EM datasets. Framed as sequence-pair classification, the models outperform classical baselines like DeepMatcher and Magellan. Their ability to handle long, unstructured records and adapt through fine-tuning underscores the strength of transformers for managing both schema and semantic heterogeneity in modern EM.

### v2

@@ Abstract
This survey offers a unified perspective on HEM by introducing a taxonomy, grounded in prior work, that distinguishes two primary categories–representation and semantic heterogeneity–and their subtypes. The taxonomy provides a systematic lens for understanding how variations in data form and meaning shape the complexity of matching tasks. We then connect this framework to the FAIR principlesFindability, Accessibility, Interoperability, and Reusability–demonstrating how they both reveal the challenges of HEM and suggest strategies for mitigating them.
@@
keywords:
Entity Matching , Entity Resolution , Data Heterogeneity
journal: Data & Knowledge Engineering
[…]
[uwoca]organization=University of Western Ontario, city=London, state=Ontario, country=Canada
[…]
[utsa]organization=University of Texas at San Antonio, city=San Antonio, state=Texas, country=USA
@@ 1 Introduction
Entity Matching (EM) has long been a fundamental component of data integration, cleaning, and analytics pipelines [23, 32, 20]. Although recent advances—especially in deep learning and AI—have accelerated progress [20, 28, 69], EM systems continue to struggle in real-world deployments. High-performing models trained on clean, benchmark datasets often fail to generalize when exposed to messy, noisy, and heterogeneous data found in practice [84, 26]. These failures are not incidental; they stem from a pervasive and under-addressed challenge: data heterogeneity. Differences in formats (e.g., dates, units), schemas (e.g., attribute names, nesting), terminology (e.g., synonyms, language), and data quality (e.g., missing or inconsistent values) introduce mismatches in structure, semantics, and quality between development and deployment settings. Such heterogeneity undermines blocking, feature extraction, and similarity computation, leading to degraded performance across the entire EM pipeline. Even recent deep learning–based methods, which perform well on standard benchmarks, suffer sharp drops in accuracy when applied across domains with varying schema or semantics [84, 68, 40].
In practice, data heterogeneity manifests in many intertwined ways, reinforcing the challenges outlined above. For example, the same product may appear as “Apple iPhone 14 (Blue)” in one source, “IPH14-BLU” in another, and only as an image with minimal text in a third—illustrating representation heterogeneity. Clinical datasets often express the same concept using terms such as “Hypertension,” “High blood pressure,” or “HTN,” revealing semantic heterogeneity. Two datasets may encode addresses differently, with one storing the full address in a single field while another splits it across multiple attributes, exemplifying structural heterogeneity. Context also varies: job titles like “Senior” or “Manager” can carry different meanings across organizations or languages, leading to contextual heterogeneity. Multilingual and multimodal environments introduce additional variation: the same city may appear as “München,” “Munich,” or “Munique,” and entities may be represented as tables, JSON records, knowledge-graph triples, or images. These diverse patterns illustrate the breadth and complexity of heterogeneity that EM systems must contend with in real-world deployments.
@@
While heterogeneity in data is well-recognized across many domains–including information retrieval [110], geospatial systems [43], the Internet of Things [4], and big data analytics [10]–it poses particularly acute and evolving challenges in EM. Historically, heterogeneity in EM has been handled under themes such as schema matching [102], duplicate record detection [32], and semantic integration [25], typically focusing on specific dimensions like attribute alignment mismatches or lexical variation. However, the modern data landscape–characterized by large-scale, semi-structured and unstructured sources from the web, data lakes, IoT devices, and enterprise systems–introduces more complex and compounded forms of heterogeneity. These include variation in data formats (e.g., JSON, XML, relational), schemas, semantics, language, granularity, and data quality. In this context, mismatches in data models and semantic assumptions severely complicate schema alignment, feature extraction, and record linkage [8]. These challenges call for EM methods that are explicitly robust to heterogeneity, and for a systematic understanding of the many forms that heterogeneity can now take in practice.
@@
To address this challenge, we argue that categorizing and systematically studying data heterogeneity is essential for advancing the design, evaluation, and deployment of EM systems. A principled taxonomy of heterogeneity is useful for organizing prior work and also enables several concrete benefits in the design and assessment of EM systems. First, it guides the development of targeted model architectures, allowing practitioners to align method design with the expected types of heterogeneity. Prior work has demonstrated that different model families excel under specific types of heterogeneity–for example, transformer-based architectures have shown robustness to semantic variation such as synonyms and abbreviations [29, 13, 120], while graph-based methods are effective at capturing structural mismatches across schemas [141, 127, 115, 65]. Building on this foundation, our paper provides new experimental evidence supporting the need for heterogeneity-aware modeling.
@@
Second, such a taxonomy enables component-level stress testing, in which researchers can evaluate how EM methods respond to controlled semantic or structural mismatch at different stages of the pipeline–from blocking to similarity computation to final classification. Third, it supports the construction of heterogeneity-aware benchmarks and perturbation-based evaluation frameworks. Finally, an explicit understanding of heterogeneity contributes to uncertainty quantification and robustness analysis, which are increasingly critical for deploying EM systems in high-stakes domains such as healthcare, scientific data integration, and finance.
@@
One of the most significant recent shifts in the EM landscape is the increasing influence of large language models (LLMs) and generative AI. These models offer new capabilities that are particularly relevant for addressing semantic heterogeneity, a core challenge in modern EM. Pretrained models such as BERT and GPT have demonstrated strong abilities to capture lexical and contextual variation through transfer learning, reducing the need for hand-crafted features or schema-specific engineering [68]. More recently, prompting and instruction tuning have enabled the use of foundation models for zero- or few-shot entity resolution [31], making it possible to generalize across domains without extensive retraining. These trends suggest that foundation models are poised to play a growing role in heterogeneity-aware EM–a theme we revisit in detail later in the survey.
The study of heterogeneity in EM also has significant implications for data governance and interoperability, particularly in the context of the FAIR principles for scientific data management–ensuring that data is Findable, Accessible, Interoperable, and Reusable [137]. Heterogeneity presents direct challenges to achieving FAIR compliance, especially when integrating records across fragmented, inconsistent, or mismatched sources. Conversely, EM methods that are explicitly designed to handle such heterogeneity can act as critical enablers of FAIRification by enhancing schema alignment, disambiguation, and record linkage. Throughout this paper, we emphasize the mutual relationship between EM and FAIR, and show how heterogeneity-aware EM systems contribute to building more trustworthy, transparent, and reusable data infrastructures.
@@
  • We present a hierarchical taxonomy of data heterogeneity in EM—adapted from established distinctions in related areas, distinguishing between representation heterogeneity (e.g., format, schema) and semantic heterogeneity (e.g., language, granularity, quality).
  • We systematically survey and categorize recent EM methods through the lens of this taxonomy, revealing how different model classes–rule-based, neural, and graph-based–address (or fail to address) specific forms of heterogeneity, and identifying patterns and gaps in current research.
@@ 2 A Framework for Classifying HEM
In this section, we first formalize the problem setting of heterogeneous entity matching (HEM) in a general and modality-agnostic way. This provides a unified foundation for describing where heterogeneity arises and how it affects the core EM task. We then introduce our taxonomy of heterogeneity types, which builds on this formalization and organizes representation and semantic heterogeneity into a coherent hierarchy.
2.1 The HEM Problem
To discuss HEM more formally, assume $E=\{e_{1},\ldots,e_{|E|}\}$ denotes the set of real-world entities in an application domain. A mention is any observable representation of an entity in $E$ within a dataset $D$, such as a relational record, a textual span (word or phrase), an image crop, or a node in a graph. For a dataset $D$, we denote its set of mentions by $\mu(D)=\{m_{1},\ldots,m_{|D|}\}$. Each mention corresponds to an underlying entity through an unknown mapping $g_{D}:\mu(D)\rightarrow E$. Given two datasets $D$ and $D^{\prime}$, the goal of EM is to determine when two mentions in the two datasets refer to the same real-world entity. Formally, EM seeks a binary matching function $f:\mu(D)\times\mu(D^{\prime})\rightarrow\{0,1\}$, where $f(m,m^{\prime})=1$ if and only if $g_{D}(m)=g_{D^{\prime}}(m^{\prime})$. When $D=D^{\prime}$, EM reduces to recovering the latent equivalence classes in $D$ induced by $g_{D}$.
Heterogeneous EM (HEM) refers to the EM task when the datasets containing the mentions exhibit heterogeneity. Between-dataset heterogeneity arises when $D$ and $D^{\prime}$ differ in how they represent, encode, or interpret information—for example, differences in schema design (attribute names, types, or structure), terminology or linguistic usage, granularity (e.g., “USA” vs. “California”), or modality (text vs. images). Within-dataset heterogeneity arises when mentions inside a single dataset $D$ vary along these same dimensions, such as mixed attribute formats, inconsistent value representations, or multilingual text within the same source. Both forms of heterogeneity complicate the decision of whether mentions refer to the same real-world entity.
We frame our discussion around Entity Matching (EM), while remaining mindful that the underlying task appears across research communities under different but closely related terminology. Common alternatives include Entity Resolution (ER), Record Linkage, Duplicate Detection, Record Matching, and Identity Resolution [32, 18]. ER is often defined as a broader pipeline involving blocking, pairwise comparison, clustering, and inconsistency resolution [61], whereas EM typically refers more narrowly to the pairwise matching or similarity assessment stage within this pipeline [69]. Related correspondence problems also arise in adjacent settings, such as Entity Linking in natural language processing [107] and Entity Alignment in knowledge graphs [117]. We adopt EM as our central term because our focus is on the matching function itself and on how representation, schema, and semantic heterogeneity affect this component across diverse data modalities and schema environments.
@@ 2.1 The HEM Problem
Similar high-level distinctions between structural (syntactic) and semantic heterogeneity have been discussed in related areas such as schema matching [101], semantic integration [24], ontology alignment [34], and federated databases [108]. However, to our knowledge, no prior work has systematically adapted these ideas to EM or used them as an organizing framework for surveying EM methods under heterogeneity. Our goal is not to introduce a new theoretical taxonomy, but to employ this framework as a practical lens for structuring, comparing, and assessing recent EM approaches across diverse forms of heterogeneity.
@@ 2.2 Representation Heterogeneity
Representation heterogeneity encompasses the structural and syntactic differences that occur when datasets use different modalities, formats, or schema designs to describe entities. These differences can disrupt every stage of the EM pipeline–from feature extraction to similarity computation–by introducing incompatibilities in how records are organized or encoded. For example, one dataset might store product data as nested JSON objects, while another uses flat CSV files. Even when datasets describe the same entities, format mismatches and inconsistent attribute organization can hinder alignment. Addressing representation heterogeneity typically requires schema matching, format normalization, or modality-specific processing techniques. We break this category into three subtypes:
  • Multimodality: Datasets may include diverse data types—such as text, images, and videos—for describing entities. For instance, e-commerce records may pair textual descriptions with product photos [58, 71]. Aligning entities across modalities requires models that can jointly embed or compare representations across heterogeneous data sources [140, 79, 16].
@@ 2.3 Semantic Heterogeneity
Early database work recognized these challenges in federated systems [108], and Semantic Web research later addressed ontology alignment [34]. Today, modern EM methods incorporate contextual modeling, domain adaptation, and knowledge resources to address semantic heterogeneity–topics we explore in depth in Section 3.
@@ 3 Review of HEM and Related Research Areas
To systematically review existing EM methods in relation to heterogeneity, we adopted a multi-step process. We first identified a set of research areas that directly intersect with heterogeneity in EM, including schema and structural heterogeneity, representation learning, deep and graph-based models, knowledge graphs and ontologies, transfer learning and domain adaptation, active and interactive learning, self-supervised and evolutionary methods, and LLMs. While our survey emphasizes recent progress in HEM, we also include influential earlier works where necessary to provide conceptual grounding and illustrate how research in these areas has evolved.
@@
We then collected and reviewed peer-reviewed articles from leading data management, AI, and ML venues that engage with heterogeneity in any of these areas. We excluded theses, posters, and papers that do not explicitly address heterogeneity. From this broader set of publications, we identified a core group of studies that directly target HEM, with particular emphasis on semantic, structural, and format heterogeneity, given their frequent overlap in practical EM settings.
3.1 Schema Heterogeneity and EM Approaches
Schema Matching and EM are distinct but related tasks. Schema matching aims to identify correspondences between attributes or structural elements of two schemas, while EM operates at the level of entity instances. Schema matching is especially relevant to schema heterogeneity as a common practical obstacle for EM: when two datasets organize or name attributes differently, some form of schema alignment is typically required before pairwise matching can be reliably performed. For this reason, schema matching is often treated as a preliminary step in traditional EM pipelines, particularly in data integration settings [102].
In modern EM, however, a growing body of work seeks to avoid explicit schema alignment altogether by developing EM models that operate directly on heterogeneous or partially aligned schemas. Below, we review EM approaches designed to handle schema or structural variation without requiring a separate schema matching stage. These methods address schema-level representation heterogeneity in our taxonomy (Figure 1).
@@ 3.1 Schema Heterogeneity and EM Approaches
Addressing this need, HERA [70] proposes a paradigm for entity resolution that bypasses schema alignment. Instead of first integrating schemas, HERA operates directly on heterogeneous records, reducing information loss from premature schema integration. It uses a compare-and-merge process to iteratively build “super records,” combining both instance-based and schema-based similarity signals. An indexing structure supports efficient candidate generation and similarity computation. Empirical results show that HERA significantly outperforms state-of-the-art methods, particularly when schema heterogeneity is high and schema mapping is lossy.
@@
Similarly, GraphER [65] avoids explicit schema alignment through a token-centric model built on Graph Convolutional Networks (GCNs). It constructs an Entity Record Graph (ER-Graph) that encodes relationships between records, attributes, and tokens. Through a two-layer GCN, GraphER jointly learns structural and semantic token embeddings, enabling fine-grained token-level comparisons without fixed attribute alignment. Evaluations on standard benchmarks show that GraphER consistently outperforms baseline models, especially in cases with diverse or sparse schemas.
Finally, Machamp [129] introduces a benchmark for Generalized Entity Matching (GEM) that further highlights these challenges. Unlike earlier EM benchmarks that assume structured data with aligned schemas, Machamp includes structured, semi-structured (e.g., JSON), and unstructured (e.g., text) data. It covers seven real-world scenarios that reflect schema mismatches, such as matching relational with semi-structured or textual data. By repurposing and transforming existing datasets, Machamp evaluates model robustness under schema and semantic variation. Experiments reveal substantial performance drops for deep models like BERT and DITTO in heterogeneous settings, underscoring the need for more schema-agnostic techniques.
@@ 3.2 Representation Learning and Semantic Embeddings
Representation learning has become foundational to modern EM [69, 141], especially with the rise of deep neural models. Unlike static feature extraction, representation learning automatically identifies latent patterns that are most predictive for matching, enabling comparisons across diverse formats, schemas, and domains. Semantic embeddings—such as Word2Vec, GloVe, and BERT—encode the contextual meaning of tokens, attributes, and records, making them particularly effective for resolving semantic heterogeneity, including synonymy, polysemy, and language variation. Most recent EM systems use some form of learned representation to handle heterogeneity in both structure and meaning.
An early transformer-based approach is EM Transformer [13], which evaluates four transformer architectures—BERT, RoBERTa, XLNet, and DistilBERT—on noisy and textual EM datasets. Framed as sequence-pair classification, these models outperform classical baselines like DeepMatcher and Magellan. Their ability to handle long, unstructured records and adapt through fine-tuning highlights the strength of transformers for managing both schema and semantic heterogeneity.
Building on this direction, DITTO [69] fine-tunes Transformer-based models such as BERT and RoBERTa for EM tasks. Treating EM as sequence-pair classification, DITTO uses contextual embeddings to capture language structure and relational cues. Its performance is further improved by (1) domain knowledge injection to highlight key tokens, (2) TF-IDF summarization to handle long records, and (3) data augmentation to enhance robustness. These design choices reduce dependency on large labeled datasets and enable generalization across schemas without requiring attribute alignment.
Extending this line of work, HierGAT [141] introduces a Hierarchical Graph Attention Transformer that models entities, attributes, and tokens through a layered graph structure. It combines self-attention and graph attention to learn multi-level contextual embeddings. The model addresses challenges such as polysemy, attribute salience, and noisy input by capturing both semantic and relational dependencies. HierGAT demonstrates strong performance on “dirty” datasets, showing robustness to data quality issues.
@@
Complementing these architectures, [144] provide an empirical comparison of twelve off-the-shelf embeddings—including FastText, SBERT, and several BERT variants—across blocking and matching tasks. Surprisingly, they find that cosine similarity over frozen embeddings often rivals fine-tuned deep models, depending on dataset properties. These insights help practitioners select efficient and effective embedding strategies for heterogeneous matching scenarios.